OpenStudy (anonymous):

Perform the indicated operation. (3 + i) + (5 + 7i) = A) -2 + 6i B) 2 + i C) 8 + 6i D) 8 + 8i

OpenStudy (raden):

open ur bracket, then combine the similar terms (3 + i) + (5 + 7i) = (3+5) + (i + 7i) = 8 + 8i